OpenStudy (eamier):

my exam question. find the length of curve log(sin x). π/4<=x<=π/3. hint:let t=cosx.

OpenStudy (anonymous):

length of curve=\[\int\limits_{\pi/4}^{\pi/3} (1+(dy/dx))^{2}\]

OpenStudy (eamier):

the formula is power 1/2 right